# AstroPix: A Pixelated HVCMOS Sensor for Space-Based Gamma-Ray Measurement

### v1

@@ 7 Conclusion and Outlook
The AstroPix design program is strong and robust as advances in the design are consistently made toward realizing a final version. AstroPix versions one through three show consistent improvements in key metrics such as power consumption and energy resolution, as shown in Table 1 however the design process is ongoing.
This paper has presented an overview of AstroPix_v3 HVCMOS design and benchtop operation. Current-voltage measurements illustrated chip properties, including high-voltage breakdown between V. An overview of energy resolution studies and results from Ref. 16 were presented, where FWHM energy resolution at keV was achieved using a medium-resistivity cm substrate. The measured m depletion of this substrate is discussed, and plans for increasing this depletion depth to the designed m were presented. Radiation testing with a cocktail of ions was performed with AstroPix_v2 which shares an identical digital bloc to AstroPix_v3, and no catastrophic events were detected. Single event functional interrupt rates were estimated to be at the order of per day in the planned AMEGO-X orbit, which is a tolerable rate.
The first space-based test of AstroPix will be the Astropix Sounding rocket Technology dEmonstration Payload (A-STEP), featuring three AstroPix_v3 quad chips to be flown on a sounding rocket in 2025 [28]. The first large-format test of AstroPix is the AMEGO-X prototype tower ComPair2 [29], which is intended to fly on a high-altitude balloon. The other AMEGO-X subsystems and operations team take heritage from the 2023 ComPair flight [30]. ComPair2 features 10 tracker layers with nearly 100 AstroPix_v3 chips per layer, which will interact with a unified trigger system to return science data. The results from this publication serve as a baseline for AstroPix_v3 operation, and will provide the basis for A-STEP design, optimization, calibration, and analysis.
At the time of writing, AstroPix_v4 has been fabricated and is undergoing preliminary testing [18]. This testing has informed the submission of AstroPix_v5 during the 2025 calendar year. Continued improvements in power consumption and dynamic range are expected with the elimination of an external fast MHz clock and addition of dynamic feedback capacitance.
@@ Acknowledgments
ALS and DV acknowledge that research was sponsored by NASA through a contract with ORAU. KK and ZM acknowledge that this material is based upon work supported by NASA under award number 80GSFC21M0002. YS’s work was supported by JSPS KAKENHI, Japan, Grant Numbers JP23K13127.

### v2

@@ 7 Conclusion and Outlook
The AstroPix design program is strong and robust as advances in the design are consistently made toward realizing a final version. AstroPix v1 through v3 show consistent improvements in key metrics such as power consumption and energy resolution, as shown in Table 1 however the design process is ongoing.
This paper has presented an overview of AstroPix_v3 HVCMOS design and benchtop operation. Current-voltage measurements illustrated chip properties, including high-voltage breakdown between V. An overview of energy resolution studies and results from Ref. 16 were presented, where FWHM energy resolution at keV was achieved using a medium-resistivity substrate. The measured m depletion of this substrate is discussed, and plans for increasing this depletion depth to the designed m were presented. Radiation testing with a cocktail of ions was performed with AstroPix_v2 which shares an identical digital bloc to AstroPix_v3, and no catastrophic events were detected. Single event functional interrupt rates were estimated to be at the order of per day in the planned AMEGO-X orbit, which is a tolerable rate.
The first space-based test of AstroPix will be the Astropix Sounding rocket Technology dEmonstration Payload (A-STEP), featuring three AstroPix_v3 quad chips to be flown on a sounding rocket in 2026 [30]. The first large-format test of AstroPix is the AMEGO-X prototype tower ComPair2 [31], which is intended to fly on a high-altitude balloon. The other AMEGO-X subsystems and operations team take heritage from the 2023 ComPair flight [32]. ComPair2 features 10 tracker layers with 380 AstroPix chips per layer, which will interact with a unified trigger system to return science data. The results from this publication serve as a baseline for AstroPix_v3 operation, and will provide the basis for A-STEP design, optimization, calibration, and analysis.
At the time of writing, AstroPix_v4 has been fabricated and is undergoing preliminary testing [18]. This testing has informed the submission of AstroPix_v5 during the 2025 calendar year. Continued improvements in power consumption and operational range are expected with the elimination of an external fast MHz clock and addition of dynamic feedback capacitance.
@@ Acknowledgments
ALS and DV acknowledge that research was sponsored by NASA through a contract with ORAU. KK and ZM acknowledge that this material is based upon work supported by NASA under award number 80GSFC21M0002. YS’s work was supported by JSPS KAKENHI Grant Numbers JP23K13127, JP25K01028.